A block of mass 'm' is places over block of mass 'M', at   its left edge. The lower block is moving with constant velocity v. The kinetic friction between the big block & the horizontal surface is uk(mu-k), & the static friction between the upper n lower block is us(mu-s). WHAT IS THE TIME TAKEN FOR BOTH THE BLOCKS TO SEPERATE FROM EACH OTHER?

Grade:11

1 Answers

vikas askiitian expert
13 years ago

U relative=v-0=v

  a relative=us.g

 v=u-at

 t=v/us.g